# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v1

@@ Abstract
Accurate hydrological understanding and water cycle prediction are crucial for addressing scientific and societal challenges associated with the management of water resources, particularly under the dynamic influence of anthropogenic climate change. Existing reviews predominantly concentrate on the development of machine learning (ML) in this field, yet there is a clear distinction between hydrology and ML as separate paradigms. Here, we introduce physics-aware ML as a transformative approach to overcome the perceived barrier and revolutionize both fields. Specifically, we present a comprehensive review of the physics-aware ML methods, building a structured community (PaML) of existing methodologies that integrate prior physical knowledge or physics-based modeling into ML. We systematically analyze these PaML methodologies with respect to four aspects: physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning. PaML facilitates ML-aided hypotheses, accelerating insights from big data and fostering scientific discoveries. We first conduct a systematic review of hydrology in PaML, including rainfall-runoff hydrological processes and hydrodynamic processes, and highlight the most promising and challenging directions for different objectives and PaML methods. Finally, a new PaML-based hydrology platform, termed HydroPML, is released as a foundation for hydrological applications. HydroPML enhances the explainability and causality of ML and lays the groundwork for the digital water cycle’s realization. The HydroPML platform is publicly available at https://github.com/HydroPML.
@@ 1 Introduction
Many of the scientific and societal challenges to understanding and preparing for environmental change rest upon our ability to understand and predict water cycle changes [1]. Process-based hydrological models play a crucial role in understanding and managing the Earth’s water resources [2] and planning for water security and managing extremes such as floods and droughts [3]. Many hydrological processes can be described as complex physical dynamics over a wide range of spatial and temporal scales, such as hydrodynamic processes and rainfall-runoff processes. Physical approaches have achieved notable success in explaining and predicting the state changes in a hydrological process [4]. However, process-based hydrology continues to present important challenges [1]. (1) A subset of process-based hydrological methods requires significant computational resources and expertise. For example, a global flood solver requires dramatic computational resources based on a traditional finite difference solver [5]. (2) Process-based hydrological models encounter limitations due to existing knowledge gaps. For example, the two-way feedback between humans and water systems is physically agnostic, but it is crucial for the water cycles [6, 7]. A new paradigm is needed to reduce knowledge gaps and explore the unknowns. (3) Process-based hydrology models frequently struggle to rapidly exploit the information in big data. For example, abundant remote sensing observations and hydrological measurements can be utilized for parameter calibration by reducing the differences between model predictions and these big data. However, most traditional process-based hydrological models rely on iterative parameter calibration, which introduces a significant level of uncertainty. Additionally, the iterative process incurs high computational costs [8].
@@
Machine learning (ML) provides efficient alternatives for learning dynamic geophysical phenomena from massive datasets [9]. Recent works have shown that ML can generate realistic short-term hydrology predictions [10] and significantly accelerate the simulation of hydrodynamic processes [11]. Despite the tremendous progress, ML is purely data-driven by nature, which presents many limitations [12, 13]. (1) The nonlinear and chaotic nature of process-based hydrology poses significant challenges for existing data-driven frameworks, which still adhere to the fundamental rules of statistical inference. (2) ML models often struggle with maintaining physical consistency. In the absence of explicit constraints, data-driven models can produce forecasts that violate the physical laws of the hydrological process, leading to implausible outcomes or plausible, yet unexplainable results. (3) Purely data-driven ML models struggle with generalization, as they face challenges in predicting untrained variables and adapting to unseen scenarios with different distributions. This is especially pronounced in process-based systems, where nonlinearity and changing system parameters contribute to distribution shifts. (4) Purely data-driven models lack interpretability and causality, which compromises the reliability of their projections when circumstances change.
@@
To our knowledge, there is no systematic summary and analysis of physics-aware ML in general. The ever-growing quantity of physics-aware ML methods makes it difficult to find the proper one for a specific problem of dynamic processes. For example, Lawal et al. [14] study and assess physics-informed neural networks (PINNs) from different researchers’ perspectives and categorize the newly proposed PINN methods into extended PINNs, hybrid PINNs, and minimized loss techniques. A review of the three neural network frameworks [17] (i.e., physics-guided neural networks, PINNs, and physics-encoded neural networks) is presented and analyzed. In addition, Goswami et al. [18] present a review of deep neural operator networks and appropriate extensions with physics-informed deep neural operators. An overview of existing physics-guided deep learning (DL) [12] is provided, and existing physics-guided DL approaches are categorized into physics-guided loss function, physics-guided architecture design, hybrid physics-DL models, and invariant and equivariant DL models. In contrast, Willard et al. [19] categorize existing methodologies into physics-guided ML models and hybrid physics-ML frameworks. However, at present, the information about physics-aware ML in these existing reviews is still limited, and these classification systems are also confusing. Beyond summarization and categorization, insightful analysis of the benefits and limitations of existing physics-aware ML methods can help researchers understand the current research state and trends of the scientific ML community. Thus, systematic analysis of physics-aware ML is crucial to the future development of physical dynamics.
@@
There is also no review of physics-aware ML specific to process-based hydrology. Existing reviews are all related to ML in the field of hydrology, or conceptual reviews of physics-aware ML in geoscience, engineering and environmental systems, or Earth system science [20]. For instance, an introductory review of ML for hydrologic sciences [21] is intended for readers new to the field of ML. Sit et al. [22] provide a comprehensive review of DL approaches used in the water industry for generation, prediction, enhancement, and classification tasks. In addition, Reichstein et al. [23] review the development of ML in the geoscientific context, highlighting DL’s potential to overcome many of the limitations of applying ML. The challenging approaches to combining ML with physical modeling are laid out. Furthermore, differentiable modeling [13] that connects varying amounts of prior knowledge to neural networks (NNs) and trains them together is proposed to offer better interpretability, generalizability, and extrapolation capability for advanced geosciences. However, many hydrological processes can be described as dynamic systems of physical equations, such as shallow water equations in hydrodynamic processes, and mass conservation equations in rainfall-runoff processes. Physics-aware ML is important for hydrology in that it integrates physical principles and domain expertise, and enables accurate predictions and extrapolation in complex hydrological systems. Thus, building a review of physics-aware ML in process-based hydrology is urgently needed for physical hydrology.
@@
There is a significant knowledge gap between physics-aware ML and process-based hydrology. Through extensive literature analysis (Fig. 1(a)), physics-aware ML is developing rapidly in the field of ML through different combinations of physics and ML. For example, partial differential equation (PDE) solutions are represented as neural networks by including the square of the PDE residual in the loss function, which was developed in 1998 [24]. In 2019, this approach was refined further and called PINNs [25], initiating a flurry of follow-up work [26]. In 2022, vanilla PINNs began to be used in the field of hydrodynamics (flood) processes [27, 28, 29]. In the ML field, in order to improve the interpretability, causality, and generalizability of ML, and rapidly capture accurate dynamic processes, dynamic physics has also been refined into physical equations, physical properties, and different physical conditions. These fine-grained physical laws are embedded in the frameworks or modules of ML in different ways for different dynamic problems, employing end-to-end training techniques. However, the current integration of physics in hydrology into ML typically follows a simple data-driven or hybrid learning approach [30], which can lead to increased computational burden and potential biases in hydrological models. Thus, further investigation and analysis of physics-aware ML and hydrology in physics-aware ML are imperative to reduce and elucidate the knowledge gap between physics-aware ML and process-based hydrology.
@@
Bearing these concerns in mind, we make an exhaustive and comprehensive review of the physics-aware ML methods. A structured community of existing methodologies that integrates prior physical knowledge or physics-based modeling into ML is built, and called PaML. As shown in Fig. 1(b), PaML approaches are categorized into four groups based on the way physics and ML are combined: (1) physical data-guided ML: a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data; (2) physics-informed ML: a widely used approach is trained from additional information obtained by enforcing the physical constraints (for example, designing loss functions (regularization)); (3) physics-embedded ML: embedding physics in the model frameworks or modules in an end-to-end manner; and (4) physics-aware hybrid learning: directly combining pure physics-based models, such as numerical methods and hydrology models, with ML models. A systematic analysis of each of these different methods is provided. Then, we conduct a systematic review of process-based hydrology in physics-aware ML, including hydrodynamic processes and rainfall-runoff hydrological processes. Next, PaML-based hydrodynamic modeling and PaML-based rainfall-runoff hydrological forecasts are classified by different objectives and PaML methods. We further highlight the most promising and challenging directions for different application categories, aiming to identify opportunities for advancement and address existing hurdles in the field. Finally, in order to reduce the knowledge gap between physics-aware ML and process-based hydrology, a new PaML-based hydrology platform, termed HydroPML, is proposed as a foundation for applications based on hydrological processes. For example, we apply HydroPML to real-time flood forecasts.
@@
Figure 1: (a) Examine the knowledge gap between physics-aware ML and hydrology in physics-aware ML from the keyword viewpoint. (b) Conceptual framework of physics-aware ML (PaML). PaML includes physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning.
@@
  • We release an open platform, termed HydroPML, to reduce the knowledge gap between physics-aware ML and process-based hydrology. In addition, we utilize HydroPML in real-time flood forecasting applications, significantly enhancing water security and fostering resilient water management.
The rest of this paper is organized as follows. Section 2 reviews state-of-the-art physics-aware ML. Section 3 presents process-based hydrology in physics-aware ML (hydroPML), and its applications in real-time flood forecasts.
@@ 2.1 Physical Data-guided Machine Learning
PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data [54]. PDgML consists of one or a combination of deep neural networks (DNN) [31], convolutional neural networks (CNNs) [32], recurrent neural networks (RNNs) [33], generative adversarial networks (GAN) [34], graph neural networks (GNN) [35], deep reinforcement learning (DRL) [36], Transfomer [37], deep operator networks [38], and physics-discovery neural networks [39]. The objectives and limitations of different physical data-guided ML are briefly discussed in Table 1.
@@ 2.1.1 Classical Physical Data-guided Neural Networks
Spatio-temporal modeling. DL has proven to be one of the most powerful machine-learning techniques for solving complex problems using big data. Most of the data generated through smart devices today are time series, and their prediction is one of the most frequent and current problems in almost all research areas [40]. The Elman network, the first RNN, incorporated the $t$ state of a hidden unit to make predictions in data sequences [41]. Long short-term memory (LSTM) [42] recurrent networks emerge to solve the vanishing gradient problem in RNNs, which uses three gates to keep long-standing relevant information and discard irrelevant information. Physics-guided LSTM [43] is designed for the extreme events and monotonic relationships of a deep learning network for rainfall-runoff simulation. Three physical mechanisms are therefore considered to constrain the LSTM network from high flow, low flow, and monotonic properties. However, existing methods are designed under limited problem settings, like predicting 48 points or fewer [44]. Recent studies have shown the potential of Transformer to increase prediction capacity. Thus, Informer [45] is proposed to enhance the prediction capacity in the long sequence time-series forecasting problem, which validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequence time-series outputs and inputs.
@@
Additionally, an army of physical data-guided spatio-temporal models [46, 47, 48] is proposed by imposing temporal and spatial inductive biases based on RNNs and CNNs. For example, ConvLSTM [49] is a type of recurrent neural network for spatio-temporal prediction that has convolutional structures in both the input-to-state and state-to-state transitions. Furthermore, inspired by the strong representation capacity of Transformer architectures (such as bidirectional encoder representations from transformers (BERT) [50] in natural language processing, vision Transformer in computer vision [51, 52]), quite a few physics-guided Transformer-based frameworks are presented. For example, a space-time Transformer for Earth system forecasting, Earthformer [53], is presented based on an efficient space-time attention block. A global data-driven weather forecasting model, FourCastNet [54], is proposed to provide accurate short- to medium-range global predictions at 0.25°$\degree$ resolution. In addition, in order to improve prediction accuracy and uncertainty, irregular space-time grid prediction problems, and other issues, a collection of classic spatio-temporal models are proposed, such as GraphCast [55] based on GNN for medium-range global weather forecasting, physically constrained GANs for improving local distributions and spatial structure of precipitation fields [56], and spatially-irregular forecasting based on geometric deep learning [57].
@@
Dynamic system modeling. Simulating multi-dimensional PDE systems with data-driven NNs has been a renewed research topic, that dates back to the last century [58]. Physical data-guided machine learning for dynamic system modeling is designed to learn a PDE model (if one exists) from a given data set and perform accurate and efficient predictions using the learned model. Development has taken place in roughly three phases.
@@
First, any early attempt [59] at the data-driven discovery of hidden physical laws compares numerical differentiations of the experimental data with analytic derivatives of candidate functions and applies symbolic regression and an evolutionary algorithm to determine the nonlinear dynamic system. Recently, motivated by the latest development of neural networks, a fully non-mechanistic method, PDE-Net [60], is designed to accurately predict the dynamics of complex systems and uncover the underlying hidden PDE models based on learning convolution kernels (filters). Further works have extended the approach with residual CNNs [61], symbolic neural networks [62], feed-forward networks [63], and deep fully convolutional networks [64]. However, these models are fundamentally limited to discretizing the input domain with a sample-inefficient grid, rendering them unable to handle temporally or spatially sparse or non-uniform observations commonly encountered in realistic applications. Recently, some convolution operators, such as Mesh-Conv [65], have been proposed to model non-uniform structured or unstructured spatial mesh.
@@
Second, neural message passing-based models such as GNN-based model [66], autoregressive solver [67], and contrastive learning-based message passing graph neural networks (MP-GNNs) [68] are related to interaction networks, where the object’s state evolves as a function of its neighboring objects, which forms dynamic relational graphs instead of grids. Neural message passing is a general framework for supervised (physical data-guided) learning on graphs [69]. In addition to the common mesh-based neural message passing-based approaches, particle-based frameworks have been attracting a lot of interest in scientific modeling for physical systems, especially those employing GNNs [70, 71, 72]. Although neural message passing-based models can achieve geometry-adaptative learning of nonlinear PDEs with arbitrary domains, these models can apply message-passing between small-scale moving and interacting objects.
@@
Third, DRL has recently been adopted in a wide range of data-driven dynamic system modeling for its ability to solve decision-making problems that were previously out of reach due to a combination of non-linearity and high dimensionality [73, 74]. For example, a DRL-based method [75] is proposed to control a coupled 2D system involving both fluid and rigid bodies, making use of a position-based reward function.

### v4

@@ Abstract
Accurate geoscientific process understanding and water cycle prediction are crucial for addressing scientific and societal challenges associated with the management of water resources. Existing reviews predominantly concentrate on the development of machine learning (ML) in this field, yet there is a clear distinction between process-based hydrology and ML as separate paradigms. Here, we introduce physics-aware ML as a transformative approach to overcome the perceived barrier and revolutionize both fields. Specifically, we present a comprehensive review of the physics-aware ML methods, building a structured community (PaML) of existing methodologies that integrate prior physical knowledge or physics-based modeling into ML. We systematically analyze these PaML methodologies with respect to four aspects: physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning. PaML facilitates ML-aided hypotheses, accelerating insights from big data and fostering scientific discoveries. We first conduct a systematic review of hydrology in PaML, including rainfall-runoff hydrological processes and hydrodynamic processes, and highlight the most promising and challenging directions for different objectives and PaML methods. Finally, a new PaML-based hydrology platform, termed HydroPML, is released as a foundation for hydrological applications. HydroPML enhances the explainability and causality of ML and lays the groundwork for the digital water cycle’s realization. The HydroPML platform is publicly available at https://hydropml.github.io/.
@@ 1 Introduction
Numerous scientific and societal challenges associated with understanding and preparing for environmental change rest upon our ability to understand and predict water cycle changes [1]. Process-based hydrological models play a crucial role in understanding and managing the Earth’s water resources [2] and planning for water security and managing extremes such as floods and droughts [3]. Many hydrological processes can be described as complex physical dynamics spanning various spatial and temporal scales, such as hydrodynamic processes and rainfall-runoff processes. Physical methods have been highly effective in elucidating and forecasting the state changes in a hydrological process [4]. However, process-based hydrology continues to present important challenges [1]. (1) A subset of process-based hydrological methods requires significant computational resources and expertise. For example, a global flood solver requires dramatic computational resources based on a traditional finite difference solver [5]. (2) Process-based hydrological models encounter limitations due to existing knowledge gaps. For example, the two-way feedback between humans and water systems is physically agnostic, but it is crucial for the water cycles [6, 7]. A new paradigm is needed to reduce knowledge gaps and explore the unknowns. (3) Process-based hydrology models frequently struggle to rapidly exploit the information in big data. For example, abundant remote sensing observations and hydrological measurements can be utilized for parameter calibration by reducing the differences between model predictions and these big data. However, most traditional process-based hydrological models rely on iterative parameter calibration, which introduces a significant level of uncertainty. Additionally, the iterative process incurs high computational costs [8].
@@
Machine learning (ML) provides efficient alternatives for learning dynamic geophysical phenomena from massive datasets [9]. Recent works have shown that ML can generate realistic short-term hydrology predictions [10] and significantly speed up the simulation of hydrodynamic processes [11]. Despite the tremendous progress, ML is purely data-driven by nature, which presents many limitations [12, 13]. (1) The nonlinear and chaotic nature of process-based hydrology poses significant challenges for existing data-driven frameworks, which still adhere to the fundamental principles of statistical inference. (2) ML models often struggle with maintaining physical consistency. In the absence of explicit constraints, data-driven models can produce forecasts that violate the physical laws of the hydrological process, leading to implausible outcomes or plausible, yet unexplainable results. (3) Purely data-driven ML models struggle with generalization, as they face challenges in predicting untrained variables and adapting to unseen scenarios with different distributions. This is especially pronounced in process-based systems, where nonlinearity and changing system parameters contribute to distribution shifts. (4) Purely data-driven models lack interpretability and causality, which compromises the reliability of their projections when circumstances change.
@@
To our knowledge, there is no systematic summary and analysis of physics-aware ML in general. The ever-growing quantity of physics-aware ML methods makes it difficult to find the proper one for a specific problem of dynamic processes. For example, Lawal et al. [14] study and assess physics-informed neural networks (PINNs) from various researchers’ perspectives, categorizing newly introduced PINN methodologies into extended PINNs, hybrid PINNs, and techniques for minimizing loss. A review of the three neural network frameworks [17] (i.e., physics-guided neural networks, PINNs, and physics-encoded neural networks) is presented and analyzed. In addition, Goswami et al. [18] present a review of deep neural operator networks and appropriate extensions with physics-informed deep neural operators. An overview of existing physics-guided deep learning (DL) [12] is provided, and existing physics-guided DL approaches are categorized into physics-guided loss function, physics-guided architecture design, hybrid physics-DL models, and invariant and equivariant DL models. In contrast, Willard et al. [19] categorize existing methodologies into physics-guided ML models and hybrid physics-ML frameworks. However, at present, the information about physics-aware ML in these existing reviews is still limited, and these classification systems are also confusing. Beyond summarization and categorization, insightful analysis of the benefits and limitations of existing physics-aware ML methods can help researchers understand the current research state and trends of the scientific ML community. Thus, systematic analysis of physics-aware ML is crucial to the future development of physical dynamics.
@@
There is also no review of physics-aware ML specific to process-based hydrology. Existing reviews are all related to ML in the field of hydrology, or conceptual reviews of physics-aware ML in geoscience, engineering and environmental systems, or Earth system science [20]. For instance, an introductory review of ML for hydrologic sciences [21] is intended for readers new to the field of ML. Sit et al. [22] offer an extensive overview of DL methods applied in the water industry, covering tasks such as generation, enhancement, prediction, and classification. In addition, Reichstein et al. [23] review the development of ML in the geoscientific context, highlighting DL’s potential to overcome many of the limitations of applying ML. The challenging approaches to combining ML with physical modeling are laid out. Furthermore, differentiable modeling [13] that connects varying amounts of prior knowledge to neural networks (NNs) and trains them together is proposed to offer better interpretability, generalizability, and extrapolation capability for advanced geosciences. However, many hydrological processes can be described as dynamic systems of physical equations, such as shallow water equations in hydrodynamic processes, and mass conservation equations in rainfall-runoff processes. Physics-aware ML is important for hydrology in that it integrates physical principles and domain expertise, and enables accurate predictions and extrapolation in complex hydrological systems. Thus, building a review of physics-aware ML in process-based hydrology is urgently needed for physical hydrology.
@@
There is a significant knowledge gap between physics-aware ML and process-based hydrology. Through extensive literature analysis (Fig. 1(a)), physics-aware ML is developing rapidly in the field of ML through different combinations of physics and ML. For example, partial differential equation (PDE) solutions are represented as NNs by including the square of the PDE residual in the loss function, which was developed in 1998 [24]. In 2019, this approach was refined further and called PINNs [25], initiating a flurry of follow-up work [26]. In 2022, vanilla PINNs began to be used in the field of hydrodynamics (flood) processes [27, 28, 29]. In the ML field, in order to improve the interpretability, causality, and generalizability of ML, and rapidly capture accurate dynamic processes, dynamic physics has also been refined into physical equations, physical properties, and different physical conditions. These fine-grained physical laws are embedded in the frameworks or modules of ML in different ways for different dynamic problems, employing end-to-end training techniques. However, the current integration of physics in hydrology into ML typically follows a simple data-driven or hybrid learning approach [30], which can lead to increased computational burden and potential biases in hydrological models. Thus, further investigation and analysis of physics-aware ML and hydrology in physics-aware ML are imperative to reduce and elucidate the knowledge gap between physics-aware ML and process-based hydrology.
@@
Bearing these concerns in mind, we make an exhaustive and comprehensive review of the physics-aware ML methods. A structured community of existing methodologies that integrates prior physical knowledge or physics-based modeling into ML is built, and called PaML. As shown in Fig. 1(b), PaML approaches are categorized into four groups based on the way physics and ML are combined: (1) physical data-guided ML: a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data; (2) physics-informed ML: a widely adopted method involves training models using supplementary information derived from enforcing physical constraints (for example, designing loss functions (regularization)); (3) physics-embedded ML: embedding physics in the model frameworks or modules in an end-to-end manner; and (4) physics-aware hybrid learning: directly combining pure physics-based models, such as numerical methods and hydrology models, with ML models. A systematic analysis of each of these different methods is provided. Then, we conduct a systematic review of process-based hydrology in physics-aware ML, including hydrodynamic processes and rainfall-runoff hydrological processes. Next, PaML-based hydrodynamic modeling and PaML-based rainfall-runoff hydrological forecasts are classified by different objectives and PaML methods. We further highlight the most promising and challenging directions for different application categories, aiming to identify opportunities for advancement and address existing hurdles in the field. Finally, in order to reduce the knowledge gap between physics-aware ML and process-based hydrology, a new PaML-based hydrology platform, termed HydroPML, is proposed as a foundation for applications based on hydrological processes. For example, we apply HydroPML to real-time flood forecasts.
@@
Figure 1: (a) Examine the knowledge gap between physics-aware ML and hydrology in physics-aware ML from the keyword viewpoint. (b) Conceptual framework of physics-aware ML (PaML) and PaML-based hydrological application highlights. PaML includes physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning.
@@
  • We release an open platform, termed HydroPML, as a foundation for applications based on hydrological processes. HydroPML bridges the gap between PaML and process-based hydrology, offering a range of hydrology applications (as exemplified in Fig. 1(b)), including but not limited to rainfall-runoff-inundation modeling and forecasting, real-time flood modeling and forecasting, and cutting-edge PaML methods to enhance water security and fostering resilient water management.
The rest of this paper is organized as follows. Section 2 reviews state-of-the-art physics-aware ML. Section 3 presents process-based hydrology in physics-aware ML (hydroPML), and its applications in real-time flood forecasts. Section 4 concludes the paper.
@@ 2.1 Physical Data-guided Machine Learning
Table 1: Proposed Physical Data-guided Machine Learning
Categories Types Examples Objective(s) Limitation(s)
Classical Physical Data-guided Neural Networks Spatio-temporal modeling LSTM [31] Three gates to keep long-standing relevant information and discard irrelevant information Limited generalization, high computational cost, limited physical interpretability
Physics-guided LSTM [32] A deep learning network for rainfall-runoff simulation
Informer [33] Long sequence time-series forecasting
ConvLSTM [34] A spatio-temporal sequence forecasting problem
Earthformer [35] A space-time Transformer for Earth system forecasting
Geometric deep learning [36] A spatially-irregular forecast method
Dynamical system modeling PDE-Net [37] A feed-forward deep network-based PDE solver Unable to handle temporally or spatially sparse or non-uniform observations
Mesh-Conv [38] A solver for modeling non-uniform structured or unstructured spatial mesh High computational cost
Neural message passing-based models [39, 40, 41, 42, 43, 44, 45] A general framework for supervised learning on graphs (arbitrary domains) Mesh-based or particle-based; message-passing between small-scale moving and interacting objects
DRL-based method [46] A coupled 2D control system Limited capabilities and robustness; unknown behavior and convergence speed in action spaces of high dimensionalities
Deep Operator Networks Deep Operator Networks (DeepONet) [47] Multiple input DeepONet [48] Employ DeepONet for realistic setups (multiple initial conditions and boundary conditions) Inaccurate approximation for complex physical dynamics
HyperDeepONet [49] Enable the learning of the DeepONet with a smaller set of parameters
Fourier neural operator (FNOs) [50] Geo-FNO [51], Domain agnostic FNO [52] Solve PDEs on arbitrary geometries Unstable, limited generalization, the basis functions in FFTs are generally frequency localised with no spatial resolution
Factorized FNO [53] Improve representation stability for the operator
Wavelet neural operatos (WNOs) [54] WNO [54] Learn the network parameters in the wavelet space Complexity, limited interpretability
Coupled multiwavelets neural operator [55] Learn coupled PDEs in the wavelet space
Graph neural operators (GNOs) [56] Graph kernel networks [56] An integral neural network that parameterizes Green’s function Difficult to train, computational cost, limited generalization
Nonlocal neural operator [57] A resolution-independent nonlocal kernel network
Physics-discovery Neural Networks Physics-discovery with unavailable prior knowledge Symbolic neural networks (SNNs) [58, 59, 60] Infer concise equations without any prior knowledge Challenging task due to complex and nonlinear systems, as well as noisy and incomplete data
AI Feynman [61, 62] Discover generalized symmetries in complex data
GNN [63] Identify the nontrivial relationship due to its exceptional inductive capability
Knowledge-agnostic discovery [6, 7] Discover the two-way feedback between human and water systems
Physics-discovery with available prior knowledge SINDy [64] Discover governing equations from data by sparse identification of nonlinear dynamical systems Limited high-fidelity noiseless measurement
PDE-FIND [65] Data-driven discovery for PDEs
PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data. PDgML consists of one or a combination of deep neural networks (DNN) [66], convolutional neural networks (CNNs) [67], recurrent neural networks (RNNs) [68], graph neural networks (GNN) [69], generative adversarial networks (GAN) [70], Transfomer [71], deep reinforcement learning (DRL) [72], deep operator networks [73], and physics-discovery neural networks [74]. The objectives and limitations of different physical data-guided ML are briefly discussed in Table 1.
@@ 2.1.1 Classical Physical Data-guided Neural Networks
Spatio-temporal modeling. DL has become a powerful tool for solving complex problems involving large datasets. Today, most of the data generated by smart devices are time series, and predicting these data is a common issue in almost every research field [75]. The Elman network, recognized as the pioneering RNN, introduce the concept of incorporating the $t$ state of a hidden unit to enhance sequence prediction [76]. To address the vanishing gradient issue inherent in RNNs, long short-term memory (LSTM) networks are developed [31]. These LSTM networks utilize three distinct gates to retain crucial information over extended periods while discarding what is deemed irrelevant. Physics-guided LSTM [32] is developed to handle extreme events and monotonic relationships in rainfall-runoff simulations using deep learning. It incorporates three physical mechanisms to regulate the LSTM network, ensuring it effectively manages high flow, low flow, and monotonic properties. However, current methods are designed within restricted problem settings, such as predicting fewer than 48 points [77]. Recent studies highlight the potential of Transformer models to improve prediction capabilities. In response, Informer [33] has been introduced to address long sequence time-series forecasting, demonstrating the Transformer-like model’s effectiveness in capturing long-range dependencies between inputs and outputs in lengthy sequences.
@@
Additionally, an army of physical data-guided spatio-temporal models [78, 79, 80] is proposed by imposing temporal and spatial inductive biases based on RNNs and CNNs. For example,ConvLSTM [34] is a recurrent neural network designed for spatio-temporal prediction, incorporating convolutional layers in both the input-to-state and state-to-state transitions. Furthermore, inspired by the strong representation capacity of Transformer architectures (such as bidirectional encoder representations from transformers (BERT) [81] in natural language processing, vision Transformer in computer vision [82, 83]), quite a few physics-guided Transformer-based frameworks are presented. For example, a space-time Transformer for Earth system forecasting, Earthformer [35], is presented based on an efficient space-time attention block. A global data-driven weather forecasting model, FourCastNet [84], is proposed to provide accurate short- to medium-range global predictions at 0.25°$\degree$ resolution. In addition, in order to improve prediction accuracy and uncertainty, irregular space-time grid prediction problems, and other issues, a collection of classic spatio-temporal models are proposed, such as GraphCast [85] based on GNN for medium-range global weather forecasting, physically constrained GANs for improving local distributions and spatial structure of precipitation fields [86], and spatially-irregular forecasting based on geometric deep learning [36].
@@
Dynamic system modeling. Simulating multi-dimensional PDE systems with data-driven NNs has been a renewed research topic, that dates back to the last century [87]. PDgML for dynamic system modeling aims to learn a PDE model (if available) from a provided dataset and make accurate and efficient predictions using the learned model. Development has taken place in roughly three phases.
@@
First, any early attempt [88] at the data-driven discovery of hidden physical laws compares numerical differentiations of the experimental data with analytic gradients of candidate functions. This approach employs symbolic regression and an evolutionary algorithm to determine the nonlinear dynamic system. Recently, motivated by the latest development of neural networks, a fully non-mechanistic method, PDE-Net [37], is designed to predict the dynamics of complex systems accurately and uncover the underlying hidden PDE models based on learning convolution kernels (filters). Further works have extended the approach with residual CNNs [89], symbolic neural networks [90], feed-forward networks [91], and deep fully convolutional networks [92]. However, these models are fundamentally constrained by their reliance on discretizing the input domain using a sample-inefficient grid. As a result, they struggle to effectively manage temporally or spatially sparse and non-uniform observations, which are frequently encountered in practical applications. Recently, some convolution operators, such as Mesh-Conv [38], have been proposed to model non-uniform structured or unstructured spatial mesh.
@@
Second, neural message passing-based models such as GNN-based model [39], autoregressive solver [40], and contrastive learning-based message passing graph neural networks (MP-GNNs) [41], are related to interaction networks where the state of an object evolves as a function of its neighboring objects, forming dynamic relational graphs instead of grids. Neural message passing is a general framework for supervised (physical data-guided) learning on graphs [42]. In addition to the common mesh-based neural message passing-based approaches, particle-based frameworks have attracted considerable interest in scientific modeling for physical systems, particularly those employing GNNs [43, 44, 45]. Although neural message passing-based models can achieve geometry-adaptive learning of nonlinear PDEs with arbitrary domains, these models can apply message-passing between small-scale moving and interacting objects.
@@
Third, DRL has been increasingly utilized in various data-driven dynamic system modeling due to its capability to address complex decision-making problems characterized by non-linearity and high dimensionality [93, 94]. For instance, a DRL-based method [46] has been developed to manage a coupled 2D system with both fluid and rigid bodies, utilizing a position-based reward function.